OpenStudy (anonymous):

y varies directly as the square of x. When x = 2, y = 40. Find x when y = 640. a. x = ±6 b. x = ±8 c. x = ±12 d. x = ±20

OpenStudy (anonymous):

ans will b B.x=±8

OpenStudy (anonymous):

thank you, if you dont mind can you explain to me how you got that?

OpenStudy (asnaseer):

"y varies directly as the square of x" implies:\[y=kx^2\]where k is some constant

OpenStudy (anonymous):

it means y = k x^2 (for some constant k) => y1/y2 = (x1/x2)^2 here y1= 40,x1 =2 and y2 = 640, x2= ?

OpenStudy (asnaseer):

you are then given: When x = 2, y = 40

OpenStudy (asnaseer):

so if you plug those values into that expression, then you can calculate the value of k.

OpenStudy (asnaseer):

\[40=k\times(2)^2=4k\implies k=10\]

OpenStudy (asnaseer):

so now you know that:\[y=10x^2\]

OpenStudy (asnaseer):

you should be able to solve from here.

OpenStudy (anonymous):

okay, i see what you did. Thanks again 640 = 10x2 so i just solve from there?

OpenStudy (asnaseer):

yes

OpenStudy (asnaseer):

what @mysha did was use a short-cut method that avoids having to calculate the value of k. I can explain that method further if you want?

OpenStudy (anonymous):

sure, i would appriciate that

OpenStudy (anonymous):

but the method was right wasn't it?

OpenStudy (asnaseer):

ok, the initial values that you were given were "When x = 2, y = 40" this can be generalised to: \(x=x_1, y=y_1\) (yes mysha your method was correct)

OpenStudy (asnaseer):

so, if we take the generalised initial values, we get:\[y_1=kx_1^2\]do you understand so far?

OpenStudy (anonymous):

@bombshellbri nw u can do it asnaseer has helped u so far alot! try it

OpenStudy (asnaseer):

here I am using \(y_1\) in place of 40 and \(x_1\) in place of 2

OpenStudy (asnaseer):

if you are not too familiar with this type of approach then we can stick to numbers and say:\[40=k\times(2)^2=4k\]

OpenStudy (asnaseer):

ok?

OpenStudy (asnaseer):

this is the same step as we did above but we do not now go on to solve for k.

OpenStudy (anonymous):

yes, i understand everything you have told me. thank you @asnaseer and you too @mysha

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

haha n.p :) credit goes to @asnaseer more ;)

OpenStudy (asnaseer):

thx @mysha :)